The upper and lower bases of a trapezoid are 13 cm and 19 cm, respectively. If the median of the trapezoid is 3x+1, find the value of x.
Answer by fractalier(6550)

The median of a trapezoid is the average of the base lengths...
thus we can write
3x + 1 = (13+19)/2 = 16
3x = 15
x = 5